OpenStudy (anonymous):

A severe storm on January 10, 1992, caused a cargo ship near the Aleutian Islands to spill 29,000 rubber ducks and other bath toys into the ocean. Eight months later, hundreds of rubber ducks began to appear along a shoreline roughly 1800 miles away. What was the approximate average speed (in mi/h and m/s) of the ocean current that carried the ducks to shore?

OpenStudy (trancenova):

Bwah ha ha, thats a funny question. Anyway laughter aside, the average speed of the rubber duckies would be: 8months = 5760hours (assuming 1mnth=30days) velocity=distance/time v=1800/5760 = 0.3125mi/h 0.3125mi/h*1h/3600s*5280f/mi*0.3048m/f = 0.1397m/s. (Could re-calculate using 5856hours; wolfram alpha said that's the amount of hours between January 10, 1992 - September 10, 1992)